In the properties of triangle if A=60 then find the value of  a /(b+c) +b/(c+a)= 

Grade:12th pass

1 Answers

Rajat
5 years ago
The question is incorrect as the value is not constant always.
Let me give a counter example: 
1)If triangle is equilateral then A= 60°
In this case a/(b+c)+b/(c+a)=1
2) if the triangle is right angled and A= 60°
B=90°, C=30° 
Then b=x, a= √3/2x , c= 1/2x
So, here a/(b+c)+b/(c+a)= 1/√3 + 2/(√3+1)= 1.309.
Please rectify the question then I will be happy to help you.!